OpenStudy (anonymous):

write the following in simplified radical form √4:(80)

OpenStudy (anonymous):

\[√4:(80) = \frac{√4}{(80)}\]

OpenStudy (anonymous):

\[√4:(80)=\frac{√4}{80}=\frac{2}{80}=\frac{1}{40}\]